Demand and supply in a market are described by the equations
Qd = 66-3P
Qs = -4+2P
(i) Solve algebraically to find equilibrium P and Q
1
Expert's answer
2018-10-11T10:24:09-0400

Equilibrium: Qd = Qs

66-3P = -4+2P

66+4 = 2P+3P

70 = 5 P

P = 70/5 = 14

Q = -4+2*14 = 24